# Charge Regulation Effect on Nanoparticles Interaction Mediated by Polyelectrolyte

### v1

@@ Abstract
The ability to precisely control surface charge using charged polymers is fundamental to many nanotechnology applications, enabling the design and fabrication of materials with tailored properties and functionalities. Here, we study the effect of charge regulation (CR) on the interaction between two nanoparticles (NPs) mediated by an oppositely charged polyelectrolyte (PE) in an electrolyte solution. To this end, we employ a hybrid CR Monte Carlo (CR-MC)/molecular dynamics (MD) simulation framework to systematically explore the effects of pH, salt concentration, and polymer chain length on NP surface charge behavior. For comparison, we also conduct MD simulations under constant charge (CC) conditions. Our results reveal that CR enhances PE adsorption onto NP surfaces compared to the CC case, where polymer bridging dominates across a wide range of NP intersurface separations. This enhanced adsorption under CR leads to a weak net repulsion driven by osmotic forces. In contrast, the CC model yields a stronger net attraction due to the bridging force. Furthermore, we find that the CR effects are more pronounced at a low salt concentration, whereas at a high salt concentration, counterion screening dominates in both CR and CC cases, resulting in similar interaction profiles. These findings highlight the importance of incorporating charge regulation in characterizing NP interactions within a complex biochemical environment, particularly in the presence of low salt concentrations.
@@
Nanoparticles (NPs) have unique properties due to their nanoscale dimensions, particularly their physicochemical properties and high surface area-to-volume ratio Deirram et al. (2019); Lin et al. (2020); Ly et al. (2024). These distinctive characteristics make them exceptionally suitable for loading materials and interacting with biological systems, leading to enhanced bioreactivity Deirram et al. (2019); Lin et al. (2020); Ly et al. (2024). Furthermore, charged polymers or polyelectrolytes (PEs) offer a versatile tool for precisely tuning the physical and chemical properties of these nanoparticles Granfeldt et al. (1991); Muthukumar (2023); Deirram et al. (2019); Mateos-Maroto et al. (2021); Ulrich et al. (2005, 2006); Carnal and Stoll (2011); Longo et al. (2013); Stornes et al. (2021); Lu et al. (2021). This is crucial for various applications, especially in the medical de la Escosura-Muñiz and Merkoçi (2012); Prakash et al. (2012); Curtis et al. (2018); Deirram et al. (2019); Lin et al. (2020); Veider et al. (2024); Ly et al. (2024) and industrial fields Karnik et al. (2005); Lu et al. (2021); Ritt et al. (2022). Interestingly, in response to their local bathing environment, both PE and NP can dynamically adjust their ionizable surface groups (consequently modulating their electrostatic interaction) Muthukumar (2023); Yuan et al. (2022); Curk et al. (2022); Curk and Luijten (2021); Bian et al. (2021); Ruixuan et al. (2023); Kumari and Podgornik (2024). Such a charge controlling mechanism is commonly called “charge regulation” (CR) and it underpins a wealth of remarkable yet complicated functional and structural motifs Lund and Jönsson (2013); Muthukumar (2023); Yuan et al. (2022); Curk et al. (2022); Curk and Luijten (2021); Bian et al. (2021); Xiong et al. (2021). Nevertheless, to take advantage of NPs in different applications, we first need a better fundamental understanding of the adverse effects of CR mechanisms on NPs interactions to predict their behavior in complex PE solutions.
@@
In the low salt concentration limit, the interaction forces between two like-charged species are primarily governed by strong repulsive forces. This repulsion arises from the overlap of diffusive double layers formed by counterions covering the charged surfaces Israelachvili (2011). A single PE in a system of two or more opposed colloids (1nm to 10 nm) mediates an attractive force by acting as a bridge Åkesson et al. (1989); Miklavic et al. (1990); Podgornik et al. (1995); Podgornik (2003). These bridging interactions arise from the fact that some of the monomers of PE are adsorbed on both surfaces, and the rest of the monomers remain in the middle, which provides an entropy-elastic force between them Podgornik et al. (1995). An attempt to describe the PE complexes onto membrane was first elucidated by Muthukumar within the framework of mean field arguments Muthukumar (1987). Based on the Edwards–deGennes self-consistent mean-field theory (SCMF) Edwards (1965); Gennes (1969) on the screened Coulomb potential interactions, Opheusden Opheusden (1988) showed short-range strong attractive bridging interactions Opheusden (1988). However, the direct surface-surface contact, which is inherently repulsive, was not included in the study Opheusden (1988). An extension of the traditional mean-field theory was developed to incorporate non-electrostatic interactions, particularly excluded volume effects between monomers Borukhov et al. (1995, 1997, 1999); Biesheuvel (2005). The SCMF was successfully applied to planar interfaces and PE adsorption on non-planar-like charged macroions Gurovitch and Sens (1999) but failed to explain the PE-mediated interaction between these small macroions. In this particular case, another elegant formulation based on a Gaussian variational ansatz Podgornik (1993) for the effective Hamiltonian of the free Podgornik et al. (1995) or grafted Podgornik (2003) PE chains was found to be more convenient Podgornik (1991, 1993). In addition, the phenomenon of “overcharging” Vagharchakian et al. (2004); Dobrynin et al. (2001); Grosberg et al. (2002); Lenz and Holm (2008) occurs in PE adsorption when multiple PEs are adsorbed on a surface in excess, resulting in charge reversal of the decorated surface. This can be understood through the concepts of correlated manner, where like-charged PE chains repel each other Dobrynin et al. (2000), and “charge fractionalization,” where the charge of a PE is not completely transferred to the surface Nguyen and Shklovskii (2002a, b). In the limit of higher salt concentration, the repulsive component of the interaction is partly due to directly screened electrostatic repulsions and partly due to PE brush repulsions between adsorbed monomers on charged surfaces Dahlgren et al. (1993). The existing various studies indicate that NP surface characteristics and conformation of PE are dramatically triggered by the properties of the PE itself (like chain length, flexibility, and charge density) and by solution pH and valency Podgornik et al. (1995); Podgornik and Ličer (2006); Ulrich et al. (2005, 2006); Carnal and Stoll (2011); Stornes et al. (2018).
@@
While advancements have been made in understanding the effect of PE on forces between charged surfaces (steric, bridging, and depletion interactions), many studies widely adopted the constant charge (CC) conditions Åkesson et al. (1989); Miklavic et al. (1990); Podgornik (1991, 1993); Podgornik et al. (1995); Podgornik and Ličer (2006); Szilagyi et al. (2014) or assume only one component (either PE or NPs) carries fixed charges Netz and Andelman (2003); Ulrich et al. (2005, 2006); Carnal and Stoll (2011); Mateos-Maroto et al. (2021); Stornes et al. (2018); Narambuena et al. (2021). This simplification overlooks the complex interplay of dynamic charges between both PE and NPs, which is crucial for accurate modeling and prediction of their behavior in solution. Attempts in this direction have so far been made only on PE adsorption on single NP Stornes et al. (2021) or flat surfaces Yuan and Tanaka (2024), suggesting that the concurrent effect of charge regulation on both PE and charged surfaces can lead to non-trivial and unexpected interaction effects that are not captured by the traditional constant charge CC model Stornes et al. (2021); Yuan and Tanaka (2024). In essence, a more holistic approach is needed, which considers the dynamic charges(acid/base groups) occurring on both the PE and the NP. In the 1970s, Ninham and Parsegian gave the first qualitative explanation of the CR effect, which couples the nonlinear Poisson-Boltzmann (PB) theory by incorporating variable surface charge densities governed by chemical equilibria for nonpolar planar surfaces in an aqueous solution Ninham and Parsegian (1971). However, the discrete nature of the ions and surface functional groups breaks down this model Ninham and Parsegian (1971), which has been shown to play a crucial role in the CR mechanism Bakhshandeh et al. (2019, 2020a, 2020b); Gomez et al. (2021). In several works, the CR formulation for surface binding sites was evaluated via the law of mass action Chan et al. (1975); Prieve and Ruckenstein (1976); von Grünberg (1999); Pericet-Camara et al. (2004) and, separately, by modifying the surface-site partition function (free energy) approach Longo et al. (2012, 2013); Adžić and Podgornik (2016); Markovich et al. (2014); Diamant and Andelman (1996); Everts et al. (2016), Frumkin–Fowler–Guggenheim isotherm model for screening effect and interaction between two macroions Ruixuan et al. (2023); Kumari and Podgornik (2024); Koopal et al. (2020). The conventional PB theory, while generally accurate for modeling simple systems like planar charged surfaces or colloids near charged surfaces faces limitations when dealing with complex systems like high-charge density scenarios, multivalent ions, and PE Muthukumar et al. (2010); Blanco et al. (2019); Stornes et al. (2021). Unfortunately, no purely theoretical approach is possible that covers a broad spectrum of biophysicochemical parameters to study NP–PE interactions. On the other hand, computer simulation techniques are more versatile and flexible than theoretical models, since they can capture CR dependence on multiple system parameters Curk and Luijten (2021); Curk et al. (2022); Yuan et al. (2022); Colla et al. (2024).
@@
In this paper, we present a comprehensive analysis of how CR influences the interaction between two NPs bearing dissociable surface groups in the presence of an oppositely charged, ionizable PE. We identify key effects of CR on the complexation behavior of the PE with the NPs in a salt solution. To model this system, we employ a recently developed hybrid charge-regulation Monte Carlo/molecular dynamics (CR-MC/MD) simulation approach Yuan et al. (2022). To the best of our knowledge, this is the first study to investigate NP–NP interactions mediated by a PE where both the NPs and the PE undergo charge regulation. By developing a more realistic understanding of such systems, this work offers critical insights for designing more effective technologies in areas ranging from biomedical applications to sustainable materials development.
@@
The outline of the paper is as follows: The details of the model underlying the simulation methodology and force calculation between various charge components are presented in Sec. II and  III, respectively. Next, in Sec. IV, the analytical aspects of our results are discussed. We delineate our final remarks in the last Sec. V
@@ II MODEL AND SIMULATION
We consider a system composed of two spherical NPs in the vicinity of an oppositely charged PE, as illustrated in Fig. 1. The system is immersed in a monovalent electrolyte solution (salt, protons, and hydroxyl ions) maintained at pH $7$. The electrostatic interactions in the system are characterized by the Bjerrum length, $\ell_{B}=q^{2}/4\pi\varepsilon\varepsilon_{0}k_{B}T$, which represents the distance at which the electrostatic interaction between two unit charges ($q$) equals the thermal energy ($k_{B}T$). We use a fixed Bjerrum length of $\ell_{B}=0.72$ nm in all simulations, corresponding to the dielectric properties of water at room temperature. Here, $\varepsilon$ and $\varepsilon_{0}$ denote the dielectric constant of the solvent and the vacuum permittivity, respectively. We simulated a cubic box with a linear size of $80\ell_{B}$ with periodic boundary conditions. A set of 256 base groups (positive sites) with dissociation constant $pK_{b}$ are uniformly distributed over each spherical shell ($3\ell_{B}$) and are firmly attached to them according to the electron distribution in the Thomson problem Wales and Ulker (2006). The PE is modeled as a bead-spring chain of $N$ weak acid groups (negative sites) with a dissociation constant $pK_{a}$ Baschnagel et al. (2000); Glotzer and Paul (2002); Kroger (2004); Peter and Kremer (2009). Here, an acid/base refers to particles that can acquire a $q=[0,\mp 1]$ charge. The NPs and PE are allowed to exchange ions with an electrolyte solution, although exhibiting overall electroneutrality. The ionizable groups undergo reactions $A^{0}\rightleftharpoons A^{-}+H^{+}$ and $B^{0}\rightleftharpoons B^{+}+OH^{-}$, respectively as developed by Curk and Lujiten Curk et al. (2022). For simplicity, we also assume that the monomers, surface sites, and solution mobile ions are of the same size ($2r=\ell_{B}$). In this method, the velocity-Verlet algorithm updates the system configurations, while CR-MC moves are employed to sample the ionization states Curk and Luijten (2021); Curk et al. (2022). To compare with the constant charge (CC) approximation, we assign fixed charges corresponding to the average ionization states observed in the CR simulations. Specifically, in the CC setup, each nanoparticle (NP) is modeled as a sphere with radius $4\ell_{B}$, bearing a constant charge equal to the average charge of an NP in the CR case. Likewise, each PE monomer is assigned a fixed charge equal to its average CR charge. The CC simulations are carried out using molecular dynamics under the same integration scheme and system parameters as the CR simulations. Counterions are included in both setups to maintain overall charge neutrality. All simulations were performed in reduced units with a time step of $0.005\tau$, using a Langevin thermostat with a damping constant of $1.0\tau$ to maintain the temperature at $T=1.0$.
@@
Figure 1: The initial configurations of our modeled systems. (a) A snapshot of the charge regulation (CR) simulation, where NPs coated with base groups ($pK_{b}=5$) are separated by an intersurface distance ($D=15\ell_{B}$). A polyelectrolyte (PE) chain composed of $N=40$ acid groups ($pK_{a}=5$) is positioned between the NPs. Red beads represent the charged monomers of the PE, while blue and white sites correspond to the NPs’ charged and neutral base groups, respectively. Green and magenta beads represent positively and negatively charged ions, respectively. (b) A snapshot of the constant charge (CC) simulation approximates the CR model. Here, NPs are assigned a fixed charge equal to the average charge on NPs in the CR simulation, and this charge is placed at the center of the NPs. A PE chain of $N=40$ monomers with fixed charges—matching the average charge on monomers in the CR simulation—is positioned between the NPs.
@@
We position the linear polyelectrolyte (PE) chain at the midpoint between the two nanoparticles (NPs), as illustrated in Fig. 1. The bonded interactions between adjacent PE monomers are modeled using a harmonic potential:
$U_{B}(r)=\frac{1}{2}K_{F}(r-r_{0})^{2}$ (1)
@@
$U_{LJ}(r_{ij})=\begin{cases}4\varepsilon_{LJ}\left[\left(\frac{\sigma}{r_{ij}-\Delta}\right)^{12}-\left(\frac{\sigma}{r_{ij}-\Delta}\right)^{6}\right],&r_{ij}\leq r_{c}^{*}\\ 0&r_{ij}>r_{c}^{*},\end{cases}$ (2)
@@
$R-R$ $2R-2r$
$R-r$ $R-r$
$r-r$ 00
$U_{coul}=\begin{cases}\frac{{\cal C}q_{i}q_{j}}{\varepsilon r_{ij}},&\text{if}\ r_{ij}\leq r_{long}^{*}\\ 0&r_{ij}>r_{long}^{*}.\end{cases}$ (3)

### v2

@@ Abstract
The ability to precisely control surface charge using charged polymers is fundamental to many nanotechnology applications, enabling the design and fabrication of materials with tailored properties and functionalities. Here, we study the effect of charge regulation (CR) on the interaction between two nanoparticles (NPs) mediated by an oppositely charged polyelectrolyte (PE) in an electrolyte solution. To this end, we employ a hybrid CR Monte Carlo / molecular dynamics simulation framework to systematically explore the effects of pH, salt concentration, and polymer chain length on NP surface charge behavior. For comparison, we also conduct molecular simulations under constant charge (CC) conditions. Our results reveal that CR enhances PE adsorption onto NP surfaces compared to the CC case, where polymer bridging dominates across a wide range of NP intersurface separations. This enhanced adsorption under CR leads to a weak net repulsion driven by osmotic forces. In contrast, the CC model yields a stronger net attraction due to the bridging force. Furthermore, we find that the CR effects are more pronounced at low salt concentration, whereas at high salt concentration, counterion screening dominates in both CR and CC cases, diminishing the CR effect. These findings highlight the importance of incorporating charge regulation in characterizing nanoparticle interactions within a complex biochemical environment, particularly in low salt concentrations.
@@
Nanoparticles (NPs) exhibit remarkable properties due to their nanoscale dimensions, primarily due to their high surface area-to-volume ratio and distinct physicochemical characteristics Deirram et al. (2019); Lin et al. (2020); Ly et al. (2024). These unique combinations make NP exceptionally suitable for loading materials and interacting with biological systems, leading to enhanced bioreactivity Deirram et al. (2019); Lin et al. (2020); Ly et al. (2024). Furthermore, charged polymers or polyelectrolytes (PEs) offer a versatile tool for precisely tuning the physical and chemical properties of these NPs Granfeldt et al. (1991); Muthukumar (2023); Deirram et al. (2019); Mateos-Maroto et al. (2021); Ulrich et al. (2005, 2006); Carnal and Stoll (2011); Longo et al. (2013); Stornes et al. (2021); Lu et al. (2021). This is crucial for various applications, especially in the medical de la Escosura-Muñiz and Merkoçi (2012); Prakash et al. (2012); Curtis et al. (2018); Deirram et al. (2019); Lin et al. (2020); Veider et al. (2024); Ly et al. (2024) and industrial fields Karnik et al. (2005); Lu et al. (2021); Ritt et al. (2022). Interestingly, in response to their local bathing environment, both PEs and NPs can dynamically adjust their ionizable surface groups (consequently modulating their electrostatic interaction) Muthukumar (2023); Yuan et al. (2022); Curk et al. (2022); Curk and Luijten (2021); Bian et al. (2021); Ruixuan et al. (2023); Kumari and Podgornik (2024). Such a charge controlling mechanism is commonly called “charge regulation” (CR), and it underpins a wealth of remarkable yet complicated functional and structural motifs Lund and Jönsson (2013); Muthukumar (2023); Yuan et al. (2022); Curk et al. (2022); Curk and Luijten (2021); Bian et al. (2021); Xiong et al. (2021). Nevertheless, to take advantage of NPs in different applications, we first need a better fundamental understanding of the role of CR mechanisms on NPs interactions to predict their behavior in complex PE solutions.
In the low salt concentration limit, the interaction forces between two like-charged species are primarily governed by strong repulsive forces. This repulsion arises from the overlap of diffusive double layers formed by counterions covering the charged surfaces Israelachvili (2011).
A single PE in a system of two or more colloids (1–10 nm) can mediate an attractive bridging interaction when the colloids carry charges opposite to that of the PE Åkesson et al. (1989); Miklavic et al. (1990); Podgornik et al. (1995); Podgornik (2003). These bridging interactions arise from the fact that some of the monomers of PE are adsorbed on both surfaces, and the rest of the monomers remain in the middle, which provides an entropy-elastic force between them Podgornik et al. (1995). An attempt to describe the PE complexes on the membrane was first elucidated by Muthukumar within the framework of mean field arguments Muthukumar (1987). Based on the Edwards–deGennes self-consistent mean-field theory (SCMF) Edwards (1965); Gennes (1969) on the screened Coulomb potential interactions, Opheusden showed short-range strong attractive bridging interactions Opheusden (1988). However, direct surface–surface contact, which is inherently repulsive, was not accounted for in the study Opheusden (1988).
@@
An extension of the traditional mean-field theory was developed to incorporate non-electrostatic interactions, particularly excluded volume effects between monomers Borukhov et al. (1995, 1997, 1999); Biesheuvel (2005). The SCMF was successfully applied to planar interfaces and PE adsorption on non-planar-like charged macroions Gurovitch and Sens (1999) but failed to explain the PE-mediated interaction between these small macroions. In this particular case, another elegant formulation based on a Gaussian variational ansatz Podgornik (1993) for the effective Hamiltonian of the free Podgornik et al. (1995) or grafted Podgornik (2003) PE chains was found to be more convenient Podgornik (1991, 1993). In addition, the phenomenon of “overcharging” Vagharchakian et al. (2004); Dobrynin et al. (2001); Grosberg et al. (2002); Lenz and Holm (2008) occurs in PE adsorption when multiple PEs are adsorbed on a surface in excess, resulting in charge reversal of the decorated surface. This can be understood through the concepts of correlated manner, where like-charged PE chains repel each other Dobrynin et al. (2000), and “charge fractionalization,” where the charge of a PE is not completely transferred to the surface Nguyen and Shklovskii (2002a, b). In the limit of higher salt concentration, the repulsive component of the interaction is partly due to directly screened electrostatic repulsions and partly due to PE brush repulsions between adsorbed monomers on charged surfaces Dahlgren et al. (1993). The existing various studies indicate that NP surface characteristics and conformation of PE are dramatically triggered by the properties of the PE itself (like chain length, flexibility, and charge density) and by solution pH and valency Podgornik et al. (1995); Podgornik and Ličer (2006); Ulrich et al. (2005, 2006); Carnal and Stoll (2011); Stornes et al. (2018).
@@
Numerous theoretical Africo et al. (2024); Cherstvy and Winkler (2011) and computational works Anhesini et al. (2025); de Carvalho et al. (2014) have provided valuable insights into PE adsorption on planar and curved surfaces. While advancements have been made in understanding the effect of PE on interaction between charged surfaces (steric, bridging, and depletion interactions), many studies widely adopted the constant charge (CC) conditions Åkesson et al. (1989); Miklavic et al. (1990); Podgornik (1991, 1993); Podgornik et al. (1995); Podgornik and Ličer (2006); Szilagyi et al. (2014) or assume only one component (either PE or NPs) carries fixed charges Netz and Andelman (2003); Ulrich et al. (2005, 2006); Carnal and Stoll (2011); Mateos-Maroto et al. (2021); Stornes et al. (2018); Narambuena et al. (2021). This simplification overlooks the complex interplay of dynamic charges between both PE and NPs, which is crucial for accurate modeling and prediction of their behavior in solution. Attempts in this direction have so far been made only on PE adsorption on a single NP Stornes et al. (2021) or flat surfaces Yuan and Tanaka (2024), suggesting that the concurrent effect of CR on both PE and charged surfaces can lead to non-trivial and unexpected interaction effects that are not captured by the traditional constant charge CC model Stornes et al. (2021); Yuan and Tanaka (2024). In essence, a more holistic approach is needed, which considers the dynamic charges (acid/base groups) occurring on both the PE and the NP.
In the 1970s, Ninham and Parsegian gave the first qualitative explanation of the CR effect, which couples the nonlinear Poisson-Boltzmann (PB) theory by incorporating variable surface charge densities governed by chemical equilibria for nonpolar planar surfaces in an aqueous solution Ninham and Parsegian (1971). However, the discrete nature of the ions and surface functional groups breaks down this model Ninham and Parsegian (1971), which has been shown to play a crucial role in the CR mechanism Bakhshandeh et al. (2019, 2020a, 2020b); Gomez et al. (2021). In several works, the CR formulation for surface binding sites was evaluated via the law of mass action Chan et al. (1975); Prieve and Ruckenstein (1976); von Grünberg (1999); Pericet-Camara et al. (2004) and, separately, by modifying the surface-site partition function (free energy) approach Longo et al. (2012, 2013); Adžić and Podgornik (2016); Markovich et al. (2014); Diamant and Andelman (1996); Everts et al. (2016), Frumkin–Fowler–Guggenheim isotherm model for screening effect and interaction between two macroions Ruixuan et al. (2023); Kumari and Podgornik (2024); Koopal et al. (2020). The conventional PB theory, while generally accurate for modeling simple systems like planar charged surfaces or colloids near charged surfaces, faces limitations when dealing with complex systems such as high-charge-density scenarios, multivalent ions, and PE Muthukumar et al. (2010); Blanco et al. (2019); Stornes et al. (2021). Unfortunately, no purely theoretical approach is possible that covers a broad spectrum of biophysicochemical parameters to study NP–PE interactions. On the other hand, computer simulation techniques are more versatile and flexible than theoretical models, since they can capture CR dependence on multiple system parameters Curk and Luijten (2021); Curk et al. (2022); Yuan et al. (2022); Colla et al. (2024).
@@
In this paper, we present a comprehensive analysis of how CR influences the interaction between two NPs bearing dissociable surface groups in the presence of an oppositely charged, ionizable PE. We identify key effects of CR on the complexation behavior of the PE with the NPs in a salt solution. To model this system, we employ a recently developed hybrid charge-regulation Monte Carlo/molecular dynamics (CR-MC/MD) simulation approach Yuan et al. (2022). To the best of our knowledge, this is the first study to investigate NP–NP interactions mediated by a PE where both the NPs and the PE undergo charge regulation. Although related studies have been conducted by the Ganeshan group, their approach applied charge regulation to the polymer while modeling the NPs (proteins) using fixed-charge patches of positive and negative ions Samanta et al. (2020); Samanta and Ganesan (2021). In contrast, our model allows all components to respond dynamically to the local chemical environment, offering a more realistic representation of the system. Experimental studies have shown that such comprehensive charge regulation is critical for accurately capturing NP–PE interactions: Huang et al. Huang et al. (2018) observed pH-dependent adsorption and force modulation between cationic PE and silica surfaces consistent with CR, while Popa et al. Popa et al. (2010) measured interaction forces between latex particles coated with poly (sodium 4-styrenesulfonate) that were quantitatively described only when boundary conditions of CR were considered. By developing a more realistic understanding of such systems, this work offers critical insights for designing more effective technologies in areas ranging from biomedical applications to sustainable materials development.
@@
The outline of the paper is as follows: The details of the model underlying the simulation methodology and force calculation between various charge components are presented in Sec. II and  III, respectively. Next, in Sec. IV, the analytical aspects of our results are discussed. We delineate our final remarks in the last Sec. V.
@@ II MODEL AND SIMULATION
Figure 1: The initial configurations of our modeled systems. (a) A snapshot of the charge regulation (CR) simulation, where nanoparticles (NPs) coated with base groups ($pK_{b}=5$) are separated by an intersurface distance ($D=15\ell_{B}$). A polyelectrolyte (PE) chain composed of $N=40$ acid groups ($pK_{a}=5$) is positioned between the NPs. Red beads represent the charged monomers of the PE, while blue and white sites correspond to the NPs charged and neutral base groups, respectively. Green and magenta beads represent positively and negatively charged ions, respectively. (b) A snapshot of the constant charge (CC) simulation approximates the CR model. Here, NPs are assigned a fixed charge equal to the average charge on NPs in the CR simulation, and this charge is placed at the center of the NPs. A PE chain of $N=40$ monomers, bearing fixed charges equal to the average monomer charge obtained in the CR simulation, is positioned between the NPs.
We consider a system composed of two spherical NPs in the vicinity of an oppositely charged PE, as illustrated in Fig. 1. The system is immersed in a monovalent electrolyte solution (salt, protons, and hydroxyl ions) maintained at pH =7=7. Electrostatic interactions in the system are characterized by the Bjerrum length, $\ell_{B}=q^{2}/4\pi\varepsilon\varepsilon_{0}k_{B}T$, which represents the distance at which the electrostatic interaction between two unit charges ($q$) equals the thermal energy ($k_{B}T$). We use a fixed Bjerrum length of $\ell_{B}=0.72$ nm in all simulations, corresponding to the dielectric properties of water at room temperature. Here, $\varepsilon$ and $\varepsilon_{0}$ denote the dielectric constant of the solvent and the vacuum permittivity, respectively. We simulated a cubic box with a linear size of $80\ell_{B}$ with periodic boundary conditions.
A set of 256 base groups (positive sites) with dissociation constant $pK_{b}$ is uniformly distributed over each spherical shell ($3\ell_{B}$) and firmly attached to them according to the electron distribution in the Thomson problem Wales and Ulker (2006). The PE is modeled as a bead-spring chain of $N$ weak acid groups (negative sites) with a dissociation constant $pK_{a}$ Baschnagel et al. (2000); Glotzer and Paul (2002); Kroger (2004); Peter and Kremer (2009). Here, an acid/base refers to particles that can acquire a $q=[0,\mp 1]$ charge. The NPs and PE are allowed to exchange ions with an electrolyte solution, although exhibiting overall electroneutrality. The ionizable groups undergo reactions $A^{0}\rightleftharpoons A^{-}+H^{+}$ and $B^{0}\rightleftharpoons B^{+}+OH^{-}$, respectively as developed by Curk and Lujiten Curk et al. (2022). For simplicity, we also assume that the monomers, surface sites, and solution mobile ions are of the same size ($2r=\ell_{B}$). In this method, the velocity-Verlet algorithm updates the system configurations, while CR-MC moves are employed to sample the ionization states Curk and Luijten (2021); Curk et al. (2022).
To compare with the constant charge (CC) approximation, we assign fixed charges corresponding to the average ionization states observed in the CR simulations. Specifically, in the CC setup, each nanoparticle (NP) is modeled as a sphere with radius $4\ell_{B}$, bearing a constant charge equal to the average charge of an NP in the CR case. Likewise, each PE monomer is assigned a fixed charge equal to its average CR charge. The CC simulations are carried out using molecular dynamics under the same integration scheme and system parameters as the CR simulations. Counterions are included in both setups to maintain overall charge neutrality. All simulations were performed in reduced units with a time step of $0.005\tau$, using a Langevin thermostat with a damping constant of $1.0\tau$ to maintain the temperature at $T=1.0$.
@@
We position the linear PE chain at the midpoint between the two NPs, as illustrated in Fig. 1. The bonded interactions between adjacent PE monomers are modeled using a harmonic potential:
$U_{B}(r)=\frac{1}{2}K_{F}(r-r_{0})^{2}~,$ (1)
@@
$U_{LJ}(r_{ij})=\begin{cases}4\varepsilon_{LJ}\left[\left(\frac{\sigma}{r_{ij}-\Delta}\right)^{12}-\left(\frac{\sigma}{r_{ij}-\Delta}\right)^{6}\right],&r_{ij}\leq r_{c}^{*}~,\\ 0~,&r_{ij}>r_{c}^{*}~,\end{cases}$ (2)
@@
$U_{coul}=\begin{cases}\frac{{\cal C}}{\varepsilon}\frac{q_{i}q_{j}}{r_{ij}},&\text{if}\ r_{ij}\leq r_{long}^{*}~,\\ 0~,&r_{ij}>r_{long}^{*}~,\end{cases}$ (3)
@@
$R\leftrightarrow R$ 2R2r2R-2r
$R\leftrightarrow r$ RrR-r
$r\leftrightarrow r$ 0